OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate-covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy’s bag weighs 1.8 pounds and it cost $1.29. Write an equation in simplest form that represents the problem described here. (Points : 1) 0.24p + 1.17 = 1.8 0.24p + 1.17 = 1.29 –0.24p + 1.602 = 1.29 –0.24p + 1.602 = 1.8

OpenStudy (anonymous):

@Algorithmic

OpenStudy (anonymous):

i think b? you

OpenStudy (anonymous):

You are correct: \[.89p + .65(1.8 − p) = 1.29\] Then we distribute and combine like terms: \[.24p + 1.17 = 1.29\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are so smart

OpenStudy (anonymous):

Here is a more in-depth explanation to anyone who may be reading this thread in the future. Let p be the amount of pretzels and g be the amout of gumdrops. Since the bag weighs \(1.8\) pounds, we know that \(p + g = 1.8\). Since it cost him $1.29 and we know the individual costs, we also know that \(.89p + .65g = 1.29\). However we still need to calculate how many gumdrops were bought, hence: \[.89p + .65(1.8 - p) = 1.29\] // \(1.8 - p\) is.... \[1.8 - p = \text{let over weight that is the gumdrops weight.}\] Hence, that is how we arrived to answer choice \(B\). Peace.